OpenStudy (anonymous):

Which expression shows how 6 • 45 can be rewritten using the distributive property? A. 6 • 40 + 6 B. 6 • 4 + 6 • 5 C. 6 • 40 + 6 • 5 D. 20 • 6 + 20 • 5

OpenStudy (anonymous):

D is the answer

jimthompson5910 (jim_thompson5910):

6 * 45 = 6 * (40 + 5) = ???

OpenStudy (anonymous):

http://www.eduplace.com/math/mhm/4/04/index.html this should help :)

OpenStudy (shamil98):

The distributive property n(a+b)

OpenStudy (shamil98):

It should be as jim showed though..

OpenStudy (anonymous):

thanks ^-^

OpenStudy (anonymous):

<(^.^<) Kirby dancing XD

OpenStudy (anonymous):

(>^.^)>

OpenStudy (anonymous):

6*45 = 6*(45)=6*(40+5) = 6*40 + 6* 5

OpenStudy (shamil98):

What he said ^

OpenStudy (anonymous):

<(^.^<) (>^.^)> <(^.^<) (>@.@>)

OpenStudy (anonymous):

I think it is D

OpenStudy (anonymous):

it is C. 6 * 45 = 6 * (40 + 5) = 6 * 40 + 6 * 5

OpenStudy (anonymous):

56 DocLav Best Response Medals 0 it is C. 6 * 45 = 6 * (40 + 5) = 6 * 40 + 6 * 5